How many grams of chlorine gas are needed to make 6.10 x10 g of a solution that is 1.90 ppm chiorine by mass? Number g CI
Answer #1

Given:-

mass of solution = 6.10 \times 106 g

concentration of chlorine solution (ppm) = 1.90

mass of chlorine gas = ?

As we know that

Concentration in ppm = mass of solute \times 106 / mass of solution

so

mass of solute =  Concentration in ppm \times mass of solution / 106

therefore

mass of chlorine gas = concentration of chlorine solution (ppm) \times mass of solution / 106

mass of chlorine gas = 1.90 \times 6.10 \times 106 g / 106

mass of chlorine gas = 1.90 \times 6.10 g

mass of chlorine gas = 11.59 g (i.e the answer)